OpenStudy (anonymous):

Solve by substation x+y=4 and x+2y=13

OpenStudy (anonymous):

\(x+y=4\) is the same as \(y=4-x\) put \(4-x\) where you see \(y\) in the second equation and get \[x+2(4-x)=13\] then solve it for \(x\)

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

or you can start with \(x=4-y\) and make the second equation \[4-y+2y=13\] and solve for \(y\) that might be even easier